PLEASE SHOW ALL WORK ! THANK YOU !!

About 30 percent of people do not recognize the bitter taster on paper impregnated with PTC. Inability to taste is due to a single autosomal recessive gene. What is the frequency of the dominant and recessive allele at this taster locus? Then what are the genotypic frequencies for this data?
Answer #1

Use Hardy-Weinberg equilibrium

p2+2pq+q2=1

p = frequency of the dominant allele in the population

q = frequency of the recessive allele in the population

p=?

q=0.30

p+q=1

p=1-q

= 1-0.30

= 0.70

Therefore, the expected frequency for dominant allele is 70%.

The expected frequency of recessive allele is 30%.

Substitute the values in the above equation:

= (0.70)2+2(0.70)(0.30)+(0.30)2

= 0.49+0.42+0.09

=1

Therefore, the genotypic frequencies of the data is 10%.